Grade 12th passGeneral Physics

The resultant of two vectors A and B is perpendicular to vector A and its magnitude is equal to half the magnitude of B. The angle between A and B is

Vedu B Das
9 Years agoGrade 12th pass

4 Answers

Vikas TU
9 Years ago
The resultant of u and v is R.
We can write the Resultant force R into vcosѲ and vsinѲ.
As per question
R = v cosѲ
 And R = v/2
 v/2 = v cosѲ
½ = cosѲ
Angle = 60 orr pi/3
Popindra Kumar
8 Years ago
The resultant of u and v is R. Here, R= v÷ 2. R= vsin∆ then sin∆=(v÷2)÷v=1÷2=sin30∆=30 then angle is 180-∆=180-30=150
Mahvish sadaf
8 Years ago
R=b/2 ( given)
Cosalpha = R/b 
=b/2/b 
1/2 
Cos alpha = 60 degree
Angle between A Nd  B = 90+ alpha
= 90+60
=150 degree 
( Bcoz R is perpendicular to A is angle between R nd A = 90 degree . The angle between B nd R  is alpha = 60  degrees .) 
  • Note :  draw a corresponding diagram to undestande better . 
  • It's a parallelogram . 
  • U find 2  triangles 
  • As angle RA = 90 the alternate angle of other triangle is also equal to 90 degree.
  • Apply cosine become we have  to relate R and B . 
  • Proud to be an IIT student
 
Navraj
8 Years ago
Since r is perpendicular to a.
Angle between a and b is π-0 
Sin0=R/B =B/2B=1/2
So 0 =30°  therefore angle between A and B is 150°